# From Many Models, One: Macroeconomic Forecasting with Reservoir Ensembles

### v1

@@ 4.2 Multi-Frequency ESNs
Ballarin et al., 2024a introduced a class of Echo State Network models designed to handle time series observed at multiple sampling frequencies. These models are collectively referred to as Multi-Frequency ESNs (MFESNs).
@@
The first category of MFESNs, known as Single-reservoir MFESNs (S-MFESNs), enables the forecasting of lower-frequency targets by mapping high-frequency reservoir states to low-frequency outputs through a state alignment scheme. In an aligned S-MFESN, the forecast is generated based on the most recent state corresponding to the reference low-frequency time index $t$. The second category, called Multi-reservoir MFESNs (M-MFESNs), introduces multiple state equations, each associated with one of the distinct sampling frequencies present in the input data. For example, quarterly and monthly variables can be incorporated jointly as regressors. By employing multiple reservoirs, each evolving at its own frequency, these models offer enhanced flexibility in capturing the dynamics of different subsets of input variables sampled at common frequencies.
@@
Following Ballarin et al., 2024a , we briefly introduce the M-MFESN framework. [^7] For the sake of brevity, we do not discuss explicitly S-MFESN models here, as they are empirically surpassed by multi-reservoir MFESNs, see our findings in Section 5. Suppose there are $Q$ groups of input time series with observations $\{Z^{(q)}_{t,{s}|{\kappa_{l}}}\}_{t,s}$, $Z^{(q)}_{t,{s}|{\kappa_{l}}}\in\mathbb{R}^{d_{q}}$, $q\in[Q]$, $t\in[T]$, $s\in\{0,\ldots,\kappa_{q}-1\}$, sampled at common frequencies $\{\kappa_{1},\ldots,\kappa_{Q}\}$, respectively. [^8] We adopt the tempo indexing of Ballarin et al., 2024a . To illustrate this indexing, suppose that high-frequency data is observed at fixed fractional sub-intervals with respect to the reference low-frequency index $t$. Each frequency is represented by an integer $\kappa_{q}$, and the fractional time index $s$ satisfies: $s=0$ indicates alignment with $t$, so $(t,{0}|{\kappa_{q}})\equiv t$; $s$ belongs to $\mathbb{Z}/\kappa_{q}$, meaning $(t,{\kappa_{q}}|{\kappa_{q}})\equiv(t+1,{0}|{\kappa_{q}})\equiv t+1$. For each input series group $q$, define a frequency-specific $D_{q}$-dimensional reservoir state
$X^{(q)}_{t,{s}|{\kappa_{l}}}=\alpha_{q}X^{(q)}_{t,{s-1}|{\kappa_{q}}}+(1-\alpha_{q})\sigma(A_{q}X^{(q)}_{t,{s-1}|{\kappa_{q}}}+C_{q}Z^{(q)}_{t,{s}|{\kappa_{q}}}+{\zeta}_{q}),$ (14)
@@
where $A_{q}$, $C_{q}$, and $\zeta_{q}$ are the group-specific reservoir state parameters, and $\alpha_{q}$ is the corresponding leak rate. To construct the forecast at the reference time $t$, the states are aligned by stacking the frequency-specific states evaluated at $s=0$, namely: $X_{t,Q}:=(X^{(1)\top}_{t,{0}|{\kappa_{1}}},\ldots,X^{(Q)\top}_{t,{0}|{\kappa_{Q}}})^{\top}\in\mathbb{R}^{D_{Q}}$, with $D_{Q}:=D_{1}+\ldots+D_{Q}$. The M-MFESN prediction equation is then given by
$Y_{t+1}$ $=\eta_{Q}+W_{Q}^{\top}X_{t,Q}+\epsilon_{t+1}=\eta_{Q}+\sum_{q=1}^{Q}W_{q}^{\top}X^{(q)}_{t,{0}|{\kappa_{q}}}+\epsilon_{t+1}.$ (15)
The readout coefficients $W_{Q}\in\mathbb{R}^{D_{Q}}$ thus linearly project $Y_{t+1}$ onto each of the $Q$ nonlinear frequency-specific states.
@@ 4.3 Ensemble (Multi-Frequency) ESNs
An important concern when working with models initialized with or fundamentally constructed with random weights is the impact that such randomness may have on model performance. A recent and growing body of work on the “lottery ticket” hypothesis (Frankle and Carbin,, 2018; Malach et al.,, 2020), argues, informally, that a key component of the empirical success of complex neural networks is “being lucky”, that is, drawing an initial configuration of weights that is favorable (Ma et al.,, 2021; Sreenivasan et al.,, 2022). Related research has focused explicitly on designing weight distributions that are guaranteed to be a priori advantageous for initialization (see, for example, Zhao et al., 2022 and Bolager et al., 2023). Echo State Networks, although proven to satisfy universal approximation and generalization guarantees (Grigoryeva and Ortega, 2018a, ; Gonon et al., 2020b, ; Gonon et al.,, 2023), also depend critically on the properties of the randomly drawn reservoir parameters. These draws cannot be fully controlled, and an unfavorable realization may degrade forecasting performance. Hence, to guard against flawed draws of $A$, $C$, and $\zeta$, we propose to construct ensembles of ESNs (EN-ESN).
Following the notation in Section 3, consider an ensemble of $K\geq 1$ distinct ESN models
$X^{(k)}_{t}$ $=\alpha_{k}X^{(k)}_{t-1}+(1-\alpha_{k})\sigma(\rho_{k}\overline{A}_{k}X^{(k)}_{t-1}+\gamma_{k}\overline{C}_{k}Z_{t}+\varsigma_{k}\overline{{\zeta}}_{k}),$ (16)
$Y_{t+1}$ $=b^{(k)}+W^{(k)\top}X^{(k)}_{t}+\epsilon^{(k)}_{t}$, (17)
with $k\in[K]$. Here, $\overline{A}_{k}$, $\overline{C}_{k}$, and $\overline{{\zeta}}_{k}$ are the $k$th (normalized) draws of the reservoir parameters, $\varphi_{k}:=(\alpha_{k},\rho_{k},\gamma_{k},\varsigma_{k})$ collects the $k$th reservoir hyperparameters, and $\theta^{(k)}:=(b^{(k)},W^{(k)\top})^{\top}$ denotes the vector of trainable readout coefficients. Each ESN’s forecast, $\widehat{Y}^{(k)}_{t+1}=b^{(k)}+W^{(k)\top}X^{(k)}_{t}$, can be combined into an ensemble prediction by means of a weighting vector $\bm{\omega}_{t+1}=(\omega_{t+1}^{(1)},\omega_{t+1}^{(2)},\ldots,\omega_{t+1}^{(K)})^{\top}$ as follows
$\widehat{Y}_{t+1}=\sum_{k=1}^{K}\omega^{(k)}_{t+1}\widehat{Y}^{(k)}_{t+1}=b_{K}+\sum_{k=1}^{K}\omega^{(k)}_{t+1}W^{(k)\top}X^{(k)}_{t},\kern 5.0pt{\rm with}\kern 5.0ptb_{K}:=\sum_{k=1}^{K}\omega^{(k)}_{t+1}\eta^{(k)}.$ (18)
We can similarly introduce ensemble MFESN (EN-MFESN) models, either with single- or multi-reservoir architectures, as a direct extension of the EN-ESN approach to the multi-frequency setting. Formally, a multi-frequency ensemble requires keeping track of two indices: $k$ for the MFESN model within the ensemble, and $q$ for the frequency group within each MFESN. For simplicity and clarity, however, we do not introduce explicit double-index notation, as our empirical applications there do not require explicit referencing of individual frequency groups of specific models within EN-MFESN combinations.
ESN Ensembles over Random Parameters (EN-RP-ESN).
@@
Since, as discussed above, the primary source of diversity across ESN or MFESN instances is the randomness of the state parameters, the first ensemble class we consider is obtained by re-drawing $A$, $C$, and $\zeta$ for each $k\in[K]$, as in (16), while keeping all other reservoir hyperparameters $\phi=(\alpha,\rho,\gamma,\varsigma)$ fixed. This type of reservoir ensemble, which we call RP-ESN ensemble (EN-RP-ESN), aims to span the family of models that differ only in terms of internal random weights. As a result, only moderate variation can be expected between members in terms of autoregressive state dynamics, input scaling, and possible activation function saturation.
@@ 5 GDP Forecasting with EN-MFESNs
In this section, we evaluate and compare the one-step-ahead forecasting performance of our proposed EN-MFESN against individual MFESN models and state-of-the-art approaches. To ensure a fair comparison, we use the same empirical experimental setup as in Ballarin et al., 2024a . We use the same data, preprocessing procedures, and benchmark specifications.
@@ 5.1 Data
The original dataset used in Ballarin et al., 2024a contains two groups of predictors: small-MD, comprising 9 predictors, and medium-MD, consisting of 33 predictors. The data is collected at different sampling frequencies, while the target variable is U.S. quarterly GDP growth. In this paper, we focus on a more challenging medium-MD data sample, while the results for the small-MD dataset can be made available by request.
The dataset is collected from various sources, including the Federal Reserve Bank of St. Louis Monthly (FRED-MD) and Quarterly (FRED-QD) Macroeconomic Databases (see McCracken and Ng, 2016 and McCracken and Ng, 2020, respectively, for more details), as well as daily series sourced from Refinitiv Datastream. The selection of predictors in the medium-MD dataset includes indicators from ten macroeconomic and financial categories (see Stock and Watson,, 1996, 2002). The preprocessing of macroeconomic indicators follows the standard guidelines outlined in McCracken and Ng, (2016, 2020). The full data sample spans a period from January 1st, 1990, to December 31st, 2019. We concentrate on the evaluation up until the Great Financial Crisis, covering the period from 1990Q1 to 2007Q4. The testing and online learning sample covers 2008Q1-2019Q4, encompassing a total of 48 forecasting rounds. [^9] For more in-depth information on specific data sources, data preprocessing and availability, we refer the reader to Table E.1 in Appendix E and Section 4.1 in Ballarin et al., 2024a .
@@ 5.2 Models and Ensembles
The first benchmark we include is the in-sample mean of U.S. GDP growth computed over the estimation and training sample. This serves as a natural baseline against which to compare mean squared forecasting errors. We also consider a linear AR(1) model, which is widely used in univariate time series analysis and macroeconomic forecasting (Stock and Watson,, 2002; Bai and Ng,, 2008). As our analysis illustrates, the empirical performance of a simple AR(1) specification can actually be superior to naïve forecast or that of poorly calibrated ensembles, making it a non-trivial reference point for evaluating our more sophisticated multi-frequency, nonlinear state-space models. Additionally, the two dynamic factor models (DFMs) employed in Ballarin et al., 2024a are also included. DFMs are widely used in empirical macroeconomic analysis as they enable the recovery of low-dimensional common components from high-dimensional panels of time series (Stock and Watson,, 2016; Doz and Fuleky,, 2020; Barigozzi and Hallin,, 2024).
@@
Two classes of ESN-based forecasting models are considered. The first set comprises single-frequency ESN models (S-MFESNs) with reservoir sizes $D=30$ and $D=120$, respectively. The sparsity degree of state matrices is set to $10/D$. The second group of models consists of multi-frequency ESNs (M-MFESNs) with separate reservoirs for monthly and daily data input series, allowing to capture the dynamics across both temporal resolutions. For each model class, two variants of MFESNs are considered: “A” and “B” types, which differ, for example, in the state hyperparameters used. Our implementation follows Ballarin et al., 2024a in defining MFESNs, namely, in choosing the architecture, reservoir sizes, activation functions, hyperparameters, time-series CV for the regularization parameter, and other relevant model-specific details. A summary of our model designs is provided in Appendix F.
@@
Figure 1: EN-RP-ESN ensembles constructed out of S-MFESN A (a), S-MFESN B (b), M-MFESN A (c), and M-MFESN B (d) models. Plots display MSFE relative to AR(1) baseline for ensemble aggregation methods (colored dashed lines), the median ensemble performance (black dotted line), specific draws of S-MFESN A/B and M-MFESN A/B from Ballarin et al., 2024a (black solid line), and empirical CDF of the relative MSFE across individual ensemble models (gray curve). Colored markers are included to distinguish overlapping lines.
@@ 5.3 Forecasting Performance
Figure 1 presents the results for EN-RP-ESN ensembles. These plots reveal substantial sensitivity to the random draw of reservoir state coefficients. Indeed, inspecting the performance of a particular (seed-based) draw of models S-MFESN A/B and M-MFESN A/B taken from Ballarin et al., 2024a , one can notice that although the sampling of both single and multi-reservoir B-type MFESN models is favorable, the S-MFESN and M-MFESN models of A type perform worse, with the latter well above the median relative forecasting error with respect to the AR(1) baseline. Simple averaging and constant Hedge follow each other closely in all panels, primarily due to the conservative choice of a small learning rate. In contrast, both FTL and AdaHedge deliver a consistent and robust significant increase in performance. It is important to notice that Follow-the-Leader (dotted red line), in particular, often converges towards a model located at the end of the left ECDF tail, indicating near-optimal performance within the ensemble. Although FTL dominates AdaHedge in all panels of Figure 1, the difference is smaller for M-MFESN than for S-MFESN models. This behavior is expected since M-MFESN models are generally more flexible and accurate than their S-MFESN counterparts.
@@
EN-RP-ESN (random parameters resampling)
@@
Table 2: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
@@
5.4 Combination Weights Dynamics
(a) S-MFESN A – AdaHedge
(b) S-MFESN A – Follow-the-Leader
(c) M-MFESN A – AdaHedge
(d) M-MFESN A – Follow-the-Leader
[…]
Our findings indicate that weight dynamics can differ significantly across different classes of MFESNs ensembles. In the upper row of Figure 3, both AdaHedge and FTL weights for the S-MFESN A ensembles change substantially over the forecasting exercise. For AdaHedge (panel (a) in Figure 3) the weight vector $\bm{\omega}_{\textnormal{{hdg}},t}$ remains very close to its equal-weight initialization until the second quarter of 2010. After that, a clearer ranking starts to emerge slowly, but ceases to stabilize. A similar pattern is visible for the Follow-the-Leader (panel (b) of Figure 3), whose weights $\bm{\omega}_{\textnormal{{FTL}},t}$ fluctuate heavily until 2012Q2, after which the leader remains unchanged. By contrast, the dynamics of M-MFESN A ensembles’ weights are more stable. Panels (c) and (d) of Figure 3 demonstrate that AdaHedge and FTL weights settle quickly and change slowly over time. For AdaHedge, the weights $\bm{\omega}_{\textnormal{{hdg}},t}$ stabilize after the end of the Great Financial Crisis downturn. Over 80% of the cumulative weight concentrates on just 5 experts, and the top-ranked expert receives over 30% of the weight as early as 2009Q1. Furthermore, the weights $\bm{\omega}_{\textnormal{{FTL}},t}$ of the FTL demonstrate quick convergence to a stable leader within less than one year from the start of the prediction exercise, signaling that the ensemble contains an M-MFESN A model which achieves consistent optimality in terms of MSFE.
It is important to emphasize that these marked differences in weight dynamics do not imply large differences in overall forecasting performance. Table 2 reports that MFESN A-type results for FTL and AdaHedge across single- and multi-reservoir architectures show much smaller MSFE gaps than, for example, across A- and B-type architectures. From an interpretability perspective, more stable weights provide clearer insight into the mechanism of the ensemble: Figure 3(d) strongly suggests an online form of model selection carried out iteratively during the forecasting exercise. An open question is hence whether there exist inherent trade-offs between ensemble “interpretability” (for example, provably simple weights dynamics for model instances from given classes of nonlinear models) and predictive performance. We leave both theoretical and empirical exploration of this question for future work.

### v2

@@ 4.2 Multi-Frequency ESNs
Ballarin et al., 2024a introduced a class of Echo State Network models designed for time series observed at multiple sampling frequencies. These models are collectively referred to as Multi-Frequency ESNs (MFESNs).
@@
The first category of MFESNs, known as Single-reservoir MFESNs (S-MFESNs), enables the forecasting of lower-frequency targets by mapping high-frequency reservoir states to low-frequency outputs through a state alignment scheme. In an aligned S-MFESN, the forecast is generated based on the most recent state corresponding to the reference low-frequency time index $t$. The second category, termed Multi-reservoir MFESNs (M-MFESNs), introduces multiple state equations, each associated with a distinct sampling frequency in the input data. For example, quarterly and monthly variables can be incorporated jointly as regressors. By employing multiple reservoirs evolving at their respective frequencies, these models flexibly capture the dynamics of different groups of input variables sampled at common frequencies.
@@
We now briefly introduce the M-MFESN framework following Ballarin et al., 2024a . For brevity, we do not discuss S-MFESN models explicitly, as they are empirically dominated by multi-reservoir MFESNs (see Section 5). Suppose there are $Q$ groups of input time series with observations $\{Z^{(q)}_{t,{s}|{\kappa_{q}}}\}_{t,s}$, where $Z^{(q)}_{t,{s}|{\kappa_{q}}}\in\mathbb{R}^{d_{q}}$, $\kappa_{q}\in\mathbb{N}$, $q\in[Q]$, $t\in[T]$, and $s\in\{0,\ldots,\kappa_{q}-1\}$, sampled at common frequencies $\{\kappa_{1},\ldots,\kappa_{Q}\}$. Using the tempo indexing of Ballarin et al., 2024a , the high-frequency observations occur at fixed fractional sub-intervals relative to the reference low-frequency index $t$. For each frequency $\kappa_{q}$, the fractional index $s=0$ indicates alignment with $t$, so that $(t,{0}|{\kappa_{q}})\equiv t$, and advancing $\kappa_{q}$ sub-intervals corresponds to one unit of low-frequency time, that is $(t,{\kappa_{q}}|{\kappa_{q}})\equiv(t+1,{0}|{\kappa_{q}})\equiv t+1$. For each input series group $q$, we define a frequency-specific $D_{q}$-dimensional reservoir state equation
$X^{(q)}_{t,{s}|{\kappa_{q}}}=\alpha_{q}X^{(q)}_{t,{s-1}|{\kappa_{q}}}+(1-\alpha_{q})\sigma(\rho_{q}\overline{A}_{q}X^{(q)}_{t,{s-1}|{\kappa_{q}}}+\gamma_{q}\overline{C}_{q}Z^{(q)}_{t,{s}|{\kappa_{q}}}+\varsigma_{q}\overline{\zeta}_{q}),$ (15)
@@
where $A_{q}$, $C_{q}$, and $\zeta_{q}$ are group-specific normalized reservoir parameters, and ${\varphi}_{q}:=(\alpha_{q},\rho_{q},\gamma_{q},\varsigma_{q})$ is the corresponding hyperparameter vector of the reservoir. To construct the forecast at the reference time $t$, the states are aligned by stacking the frequency-specific states evaluated at $s=0$, that is, $X_{t,Q}:=(X^{(1)\top}_{t,{0}|{\kappa_{1}}},\ldots,X^{(Q)\top}_{t,{0}|{\kappa_{Q}}})^{\top}\in\mathbb{R}^{D_{Q}}$, $D_{Q}:=\sum_{q=1}^{Q}D_{q}$. The M-MFESN prediction equation is given by
$Y_{t+1}$ $=b_{Q}+W_{Q}^{\top}X_{t,Q}+\epsilon_{t+1}=b_{Q}+\sum_{q=1}^{Q}W_{q}^{\top}X^{(q)}_{t,{0}|{\kappa_{q}}}+\epsilon_{t+1}.$ (16)
The readout $W_{Q}$ defines a linear map from the nonlinear, frequency-specific states to $Y_{t+1}$.
@@ 4.3 Ensemble (Multi-Frequency) ESNs
An important concern when working with models initialized or inherently constructed with random weights is the impact that such randomness may have on model performance. A recent and growing body of work on the “lottery ticket” hypothesis (Frankle and Carbin,, 2018; Malach et al.,, 2020), argues, informally, that a key component of the empirical success of complex neural networks is “being lucky”, that is, drawing an initial configuration of weights that is favorable (Ma et al.,, 2021; Sreenivasan et al.,, 2022). Related research has focused explicitly on designing weight distributions that are guaranteed to be a priori advantageous for initialization (see, e.g., Zhao et al., 2022 and Bolager et al., 2023). Echo State Networks, although proven to satisfy universal approximation and generalization guarantees (Grigoryeva and Ortega, 2018a, ; Gonon et al., 2020b, ; Gonon et al.,, 2023), also depend critically on the properties of randomly drawn reservoir parameters. Once used for forecasting such systems may offer very different forecasting accuracy. It is hence suggestive to construct ensembles of ESNs (EN-ESN). In our multifrequency setting we directly extend this idea to ensembles of MFESNs (EN-MFESN), with single- or multi-reservoir architectures.
Following the notation introduced in Section 4.2, consider an ensemble of $K\geq 1$ distinct MFESN models-experts of multiple reservoirs (M-MFESNs), each used to construct the corresponding expert prediction out of $Q$ groups of input series sampled at distinct frequencies as $\widehat{Y}^{(k)}_{t+1}=\widehat{b}^{(k)}_{Q,\lambda}+\widehat{W}^{(k)\top}_{Q}X^{(k)}_{t,Q}$, $k\in[K]$, with $\widehat{W}^{(k)\top}_{Q}$ and $\widehat{b}^{(k)}_{Q,\lambda}$ estimated as in (13)-(14). The ensemble forecast is constructed according to (2) as
$\widehat{Y}_{t+1}=\sum_{k=1}^{K}\omega^{(k)}_{t}\widehat{Y}^{(k)}_{t+1}=\widehat{b}_{K}+\sum_{k=1}^{K}\omega^{(k)}_{t}\widehat{W}^{(k)\top}_{Q}X^{(k)}_{t,Q},\kern 5.0pt{\rm with}\kern 5.0pt\widehat{b}_{K}:=\sum_{k=1}^{K}\omega^{(k)}_{t}\widehat{b}^{(k)}_{Q,\lambda}$. (17)
ESN Ensembles over Random Parameters (EN-MFESN-RP).
@@
Since, as discussed above, the primary source of diversity MFESN instances is the randomness of the state parameters, the first ensemble class we consider is obtained by re-drawing the state parameters for each $k\in[K]$, as in (15), while keeping all other reservoir hyperparameters $\varphi_{q}=(\alpha_{q},\rho_{q},\gamma_{q},\varsigma_{q})$ fixed (for single- and multi-reservoir cases). This type of reservoir ensemble, which we call MFESN-RP ensemble (EN-MFESN-RP), aims to span the family of models that differ only in terms of the sampled state parameters. As a result, only moderate variation can be expected between members in terms of autoregressive state dynamics, input scaling, and possible activation function saturation.
@@ 5 GDP Forecasting with EN-MFESNs
In this section, we evaluate and compare the one-step-ahead forecasting performance of our proposed EN-MFESN against individual MFESN models and state-of-the-art approaches. To ensure a fair comparison, we use the same empirical experimental setup as in Ballarin et al., 2024a : We use the same data, preprocessing procedures, and benchmark specifications.
@@ 5.1 Data
We conduct our analysis for the original dataset used in Ballarin et al., 2024a , which contains two groups of predictors: small-MD, comprising 9 predictors, and medium-MD, consisting of 33 predictors. The data is collected at different sampling frequencies, while the target variable is U.S. quarterly GDP growth. In this paper, we focus on a more challenging medium-MD data sample, while the results for the small-MD dataset can be made available by request. For more in-depth information on specific data sources, data preprocessing and availability, we refer the reader to Appendix F and Section 4.1 in Ballarin et al., 2024a .
@@ 5.2 Models and Ensembles
The first benchmark we consider is the in-sample mean of U.S. GDP growth computed over the estimation sample, which serves a natural baseline for comparing mean squared forecasting errors. We also include an AR(1) model, standard in univariate time series analysis and macroeconomic forecasting (Stock and Watson,, 2002; Bai and Ng,, 2008). Our analysis demonstrates that AR(1) can outperform naïve forecast and poorly calibrated ensembles, making it a non-trivial reference for evaluating multi-frequency nonlinear state-space models. Additionally, the two dynamic factor models (DFMs) from Ballarin et al., 2024a are also included. DFMs are widely applied in macroeconomic analysis to recover low-dimensional common components from high-dimensional time series (Stock and Watson,, 2016; Doz and Fuleky,, 2020; Barigozzi and Hallin,, 2024).
@@
Two classes of ESN-based forecasting models are considered. The first set comprises single-frequency ESN models (S-MFESNs) with reservoir sizes $D=30$ and $D=120$, respectively. The sparsity degree of state matrices is set to $10/D$. The second group of models includes multi-frequency ESNs (M-MFESNs) with separate reservoirs for monthly and daily data input series, allowing to capture the dynamics across both temporal resolutions. For each model class, two variants of MFESNs are considered: “A” and “B” types, which differ, e.g., in the state hyperparameters used. Our implementation of MFESNs follows Ballarin et al., 2024a in the choice of architecture, reservoir sizes, activation functions, hyperparameters, time-series cross validation for the regularization parameter, and other relevant model-specific details. A summary of our model designs is provided in Appendix I.
@@
Figure 1: EN-MFESN-RP ensembles constructed out of S-MFESN A (a), S-MFESN B (b), M-MFESN A (c), and M-MFESN B (d) models. Plots display MSFE relative to AR(1) baseline for ensemble aggregation methods (colored dashed lines), the median ensemble performance (black dotted line), specific draws of S-MFESN A/B and M-MFESN A/B from Ballarin et al., 2024a (black solid line), and empirical CDF of the relative MSFE across individual ensemble models (gray curve). Colored markers are included to distinguish overlapping lines.
@@ 5.3 Forecasting Performance
Figure 1 presents the results for EN-MFESN-RP ensembles and reveals substantial sensitivity to the random draw of reservoir state coefficients. In particular, individual S-MFESN A/B and M-MFESN A/B model draws taken from Ballarin et al., 2024a display heterogeneous performance: While B-type single- and multi-reservoir MFESNs perform favorably, A-type models often exhibit higher forecasting errors, in some cases exceeding the median relative MSE with respect to the AR(1) benchmark. Simple averaging and constant Hedge follow each other closely across panels, primarily due to the choice of a small learning rate. In contrast, both FTL and AdaHedge deliver consistent, sizable, and robust performance gains. It is important to notice that Follow-the-Leader often to models in the extreme left tail of the ECDF, indicating near-optimal performance within the ensemble. Although FTL dominates AdaHedge throughout, the difference is higher for S-MFESN models than for M-MFESNs, consistent with their greater flexibility and baseline accuracy of the latter.
@@
EN-MFESN-RP (random parameters resampling)
@@
Table 2: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing combinations are displayed in bold.